OpenStudy (girlstudy):

A car is initially moving at 35 km/h along a straight highway. To pass another car, it speeds up to 135 km/h in 10.5 seconds at a constant acceleration. (a) How large was the acceleration in m/s^2 (b) How large was the acceleration, in units of g= 9.80 m/s^2

OpenStudy (austin.l):

Well, first we need to translate km/h to m/s. To do that, we need to know dimensional analysis. Are you familiar with it?

OpenStudy (girlstudy):

Not really

OpenStudy (austin.l):

Okay, so how many meters are in a kilometer? How many seconds in a minute and how many minutes in an hour?

OpenStudy (girlstudy):

1 kilometer = 1000 meters 60secs in a minute 60 mins in an hour

OpenStudy (austin.l):

Okay, so \[\frac{35km}{1hr}\times\frac{1000m}{1km}\times\frac{1hr}{60min}\times\frac{1min}{60sec}=?\]

OpenStudy (girlstudy):

35000/3600 ?

OpenStudy (girlstudy):

But isnt there a formula we need to follow? like f=MA ? or something

OpenStudy (austin.l):

Which would be 9.722m/s and doing the same for 135km/h would be 37.5m/s The formula for acceleration is this: \[a=\frac{\Delta v}{t}\] Which is change in velocity over time. Our change in velocity would be 27.778 Our time would be 10.5seconds so, \[a=\frac{27.778m/s}{10.5s}\] which would be 2.6455m/s^2

OpenStudy (austin.l):

and for part b, just divide that number by 9.80m/s^2

OpenStudy (girlstudy):

ohh i see thank you so much!

OpenStudy (austin.l):

No problemo!